Jim can fill a pool using buckets of water in 30 minutes and Sue can do the same job in 45 minutes, Tony can do the same job in 90 minutes. How long would it take them to do it together.
Answer by Alan3354(69443)

Jim can fill a pool using buckets of water in 30 minutes and Sue can do the same job in 45 minutes, Tony can do the same job in 90 minutes. How long would it take them to do it together.
------------------
For this type of problem, you have to use the reciprocals.
Jim does 1/30 of the pool per minute.
Sue does 1/45
Tony does 1/90
Add those.
1/30 + 1/45 + 1/90 = 6/90 pools per minute = 1/15
--> 15 minutes per pool